# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

### v1

@@ Abstract
Privacy is a central challenge for systems that learn from sensitive data sets, especially when a system’s outputs must be continuously updated to reflect changing data. We consider the achievable error for the differentially private continual release of a basic statistic—the number of distinct items—in a stream where items may be both inserted and deleted (the turnstile model). With only insertions, existing algorithms have additive error just polylogarithmic in the length of the stream $T$. We uncover a much richer landscape in the turnstile model, even without considering memory restrictions. We show that any differentially private mechanism that handles insertions and deletions has worst-case additive error at least $T^{1/4}$ even under a relatively weak, event-level privacy definition. Then, we identify a property of the input stream, its maximum flippancy, that is low for natural data streams and for which one can give tight parameterized error guarantees. Specifically, the maximum flippancy is the largest number of times the count of a single item changes from a positive number to zero over the course of the stream. We present an item-level differentially private mechanism that, for all turnstile streams with maximum flippancy $w$, continually outputs the number of distinct elements with an $O(\sqrt{w}\cdot\mathsf{poly}\log T)$ additive error, without requiring prior knowledge of $w$. This is the best achievable error bound that depends only on $w$, for a large range of values of $w$. When $w$ is small, our mechanism provides similar error guarantees to the polylogarithmic in $T$ guarantees in the insertion-only setting, bypassing the hardness in the turnstile model.
@@ 1 Introduction
Machine learning algorithms are frequently run on sensitive data. In this context, a central challenge is to protect the privacy of individuals whose information is contained in the training set. Differential privacy [23] provides a rigorous framework for the design and analysis of algorithms that publish aggregate statistics, such as parameters of machine learning models, while preserving privacy. In this work, we focus on the model of differential privacy interchangeably called continual observation and continual release that was introduced by  Dwork et al. [26] and Chan et al. [14] to study privacy in settings when both the data and the published statics are constantly updated. One of the most fundamental statistics about a data stream is the number of distinct elements it contains (see, e.g., the book by Leskovec et al. [46]). The problem of counting distinct elements has been widely studied, starting with the work of Flajolet and Martin [33], and has numerous applications [2, 31, 48, 41, 4], including monitoring traffic on websites, the number of patients in a country’s hospitals, and the number of customers in a store. Algorithms for this problem are used as basic building blocks in more complicated data analyses.
We investigate the problem of privately counting the number of distinct elements under continual observation in the turnstile model, which allows both element insertions and deletions. In the continual release model, a data collector receives a sensitive dataset as a stream of inputs and produces, after receiving each input, an output that is accurate for all the inputs received so far. The input stream is denoted $x$ and its length (also called the time horizon) is denoted $T$. The elements come from a universe $\mathcal{U}$. Each entry in the stream represents an insertion (denoted by $+u$) or a deletion (denoted by $-u$) of some element $u\in\mathcal{U}$ or, alternatively, a no-op (denoted by $\bot$), representing that no update occurred in the current time step.
@@
Definition 1.1 (Existence vector and $\mathsf{CountDistinct}$).
Fix a universe $\mathcal{U}$ and a time horizon $T\in\mathbb{N}$. For an element $u\in\mathcal{U}$ and a stream $x\in\mathcal{U}_{\pm}^{T}$, the existence vector $f_{u}(x)\in\{0,1\}^{T}$ is an indicator vector that tracks the existence of element $u$ in $x$: specifically, for each $t\in[T]$, the value $f_{u}(x)[t]=1$ if and only if there are strictly more insertions than deletions of element $u$ in $x[1:t].$ The function $\mathsf{CountDistinct}{}:\mathcal{U}_{\pm}^{T}\to\mathbb{N}^{T}$ returns a vector of the same length as its input, where $\mathsf{CountDistinct}(x)[t]=\sum_{u\in\mathcal{U}}f_{u}(x)[t]$ for all $t\in[T]$.
The focus of our investigation is the best achievable error in the continual release model for a given time horizon $T$ and privacy parameters. We study the worst-case (over all input streams and time steps $t$) additive error of privately approximating the distinct element counts under continual release. Given an answer vector $a\in\mathbb{R}^{T}$, the error of this vector with respect to the desired function value $f(x)\in\mathbb{R}^{T}$ computed on dataset $x$ is defined as $\text{\sf ERR}_{f}(x,a)=\|f(x)-a\|_{\infty}.$ A mechanism in the continual release model is $\alpha$-accurate if it outputs a vector of answers $a$ with error $\text{\sf ERR}_{\mathsf{CountDistinct}}(x,a)\leq\alpha$ with probability 0.99.
Next, we discuss privacy. Originally, differential privacy [23] was defined in a setting where a data collector outputs the desired information about an entire dataset all at once. We call this the batch model, to contrast it with continual release. In the batch model, two datasets are called neighbors if they differ in the data of one individual. There are two natural ways to adapt this definition to the continual release model [26, 14], depending on the desired privacy guarantees.
@@
Let $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$ be two streams of length $T$. Streams $x$ and $x^{\prime}$ are event-neighbors if $x^{\prime}$ can be obtained from $x$ by replacing one entry of $x$ with $\perp$, or vice-versa. Streams $x$ and $x^{\prime}$ are item-neighbors if $x^{\prime}$ can be obtained from $x$ by replacing with $\perp$ a subset of the entries of $x$ pertaining to element $u$, for some $u\in\mathcal{U}$, or vice-versa.
Differential privacy can be defined with respect to any notion of neighboring datasets. There are two privacy parameters: $\varepsilon>0$ and $\delta\in[0,1)$. The case when $\delta=0$ is referred to as pure DP, and the general case as approximate DP. An algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private (DP) if for all pairs of neighboring datasets $x,x^{\prime}$ and all events $S$ in the output space of $\mathcal{A}$,
@@
For event-neighboring (respectively, item-neighboring) streams $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$, we say that $\mathcal{A}$ is $(\varepsilon,\delta)$-event-level-DP (respectively, item-level-DP). Observe that item-level DP imposes a more stringent requirement than event-level, since it guards against much larger changes in the input stream. To contrast with the batch setting, we refer to continual release algorithms as mechanisms.
In the batch setting, one can satisfy differential privacy with expected error $O(1/\varepsilon)$ since, for any particular $t$, the function $\mathsf{CountDistinct}(x)[t]$ has sensitivity 1—regardless of whether deletions are allowed and whether we consider event-level or item-level privacy. Privacy is more challenging in the continual release setting, where we aim to release a sequence of estimates, one for each time $t$, and we require that the privacy guarantee hold for the entire sequence of outputs. Prior work on privately estimating distinct elements in this setting considered the insertion-only model, exclusively: Bolot et al. [9] show that one can get a sequence of estimates, all of which are within additive error $poly(\log T)/\varepsilon$. Their result holds for both item-level and event-level privacy (which are essentially equivalent when considering only insertions). Follow-up work generalized their mechanism but, again, considered only insertions [36, 30]
We uncover a much richer landscape in the turnstile model, even without considering memory restrictions. We show that any differentially private mechanism that handles insertions and deletions has worst-case additive error at least $T^{1/4}$ even under event-level privacy, the weaker of the two privacy notions. To overcome this lower bound, we identify a property of the input stream, its maximum flippancy, that is low for natural data streams and for which one can give tight parameterized error guarantees. To define flippancy, recall the definition of the existence vector from Definition 1.1.
@@
Given a stream $x$ of length $T$ and an element $u\in\mathcal{U}$, the flippancy of $u$ in $x$, denoted by $\mathsf{flip}(u,x)$, is the number of pairs of adjacent entries in the existence vector $f_{u}(x)$ with different values. That is, $\mathsf{flip}(u,x)=|\{j\in[T-1]:f_{u}(x)[j]\neq f_{u}(x)[j+1]\}|.$ The maximum flippancy of a stream $x$, denoted $w_{x}$, equals $\max_{u\in\mathcal{U}}\mathsf{flip}(u,x)$.
In other words, the maximum flippancy is the largest number of times the contribution of a single item to the distinct element count changes over the course of the stream. We design item-level private mechanisms whose error scales with the maximum flippancy of the stream, even though the maximum flippancy is not an input to the mechanism. We show matching lower bounds that hold in all regimes for item-level privacy. For a large range of the flippancy parameter, we also show a matching lower bound for event-level privacy, using a different argument. This leaves a range with an intriguing gap between item-level and event-level bounds.
@@ 1.1 Our results
Our results are summarized in Table 1.
Bounds Item-Level Privacy Event-Level Privacy
@@
                    $\tilde{O}\left(\mathsf{min}\Big{(}\frac{\sqrt{w}\ (\operatorname{polylog}\ T)\sqrt{\log 1/\delta}}{\varepsilon},\frac{\left(T\log 1/\delta\right)^{1/3}}{\varepsilon^{2/3}},T\Big{)}\right)$ (Thm. 1.4)
@@
$\tilde{\Omega}\Big{(}\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/3}}{\varepsilon^{2/3}},T\Big{)}\Big{)}$ (Thm. 1.6) $\Omega\left(\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}},T\Big{)}\right)$ (Thm. 1.5)
Table 1: Summary of our results: bounds on the worst-case additive error for $\mathsf{CountDistinct}$ under event-level and item-level $(\varepsilon,\delta)$-differential privacy. All results apply for streams of length $T$ and maximum flippancy at most $w$ when $\varepsilon\leq 1$ and $\delta=o(\frac{\varepsilon}{T})$.
Our first result is a mechanism for privately approximating $\mathsf{CountDistinct}$ for turnstile streams. For streams of length $T$ with maximum flippancy $w_{x}$, this mechanism achieves item-level differential privacy with error $\operatorname{O}\left(\min(\sqrt{w_{x}}\mdot\operatorname{polylog}T,T^{1/3}\right))$, without knowing the maximum flippancy upfront. Since this mechanism is item-level-DP, it is also event-level-DP with the same parameters. The error it achieves is the best possible in terms of dependence only on $w$ for item-level DP, and this error is nearly tight for event-level DP. When $w$ is small, as is the case for many natural streams, our mechanism has error $O(\operatorname{polylog}\ T)$, similar to mechanisms for the insertion-only setting.
@@
Fix $\varepsilon,\delta\in(0,1]$ and sufficiently large $T\in\mathbb{N}$. Then, there exists an $(\varepsilon,\delta)$-item-level-DP mechanism for $\mathsf{CountDistinct}$ that, for all turnstile streams $x$ of length $T$, is $\alpha$-accurate where
@@
Theorem 1.4 can be easily extended to $\varepsilon$ bounded by any constant larger than $1$. We fixed the bound to be $1$ to simplify the presentation. Our mechanism has polynomial time and space complexity in the input parameters, although it does not achieve the typically sublinear space guarantees of streaming algorithms. We focus on optimizing the error guarantees as opposed to space complexity, since the error cannot be improved by procuring more computational resources, e.g., buying more memory.
Our lower bounds for $\mathsf{CountDistinct}$ for turnstile streams are parameterized by the maximum flippancy $w$ of the stream. For event-level DP, our lower bound matches the error guarantee of our $\mathsf{CountDistinct}$ mechanism for a large range of values of $w$, namely for all $w\leq T^{1/2}$ and $w\geq T^{2/3}$. The best achievable error for $w\in(T^{1/2},T^{2/3})$ for event-level DP remains an open question.
@@
Let $\varepsilon\in(0,1]$, $\delta=o\left(\frac{\varepsilon}{T}\right)$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists an $(\varepsilon,\delta)$-event-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then
@@
For item-level DP, our lower bound on the error matches our upper bound for all regimes of $w$ up to polylogarithmic factors.
@@
Let $\varepsilon\in(0,1]$, $\delta=o\left(\frac{\varepsilon}{T}\right)$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists an $(\varepsilon,\delta)$-item-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then
$\alpha=\tilde{\Omega}\Big{(}\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/3}}{\varepsilon^{2/3}},T)\Big{)}\Big{)}\text{ for approximate DP and }\alpha=\Omega\Big{(}\mathsf{min}\Big{(}\frac{w}{\varepsilon},\sqrt{\frac{T}{\varepsilon}},T\Big{)}\Big{)}\text{ when }\delta=0.$
All our lower bounds also hold in the strict turnstile model, where element counts never go below 00, and even in the model where each element can be added only when it is absent and deleted only when it is present (as is the case, for example, with the “like” counts on social media websites). The lower bounds apply even to offline mechanisms that receive the entire input stream before producing output; they do not rely on the mechanism’s uncertainty about what comes later in the stream.

### v3

@@ Abstract
Privacy is a central challenge for systems that learn from sensitive data sets, especially when a system’s outputs must be continuously updated to reflect changing data. We consider the achievable error for differentially private continual release of a basic statistic—the number of distinct items—in a stream where items may be both inserted and deleted (the turnstile model). With only insertions, existing algorithms have additive error just polylogarithmic in the length of the stream $T$. We uncover a much richer landscape in the turnstile model, even without considering memory restrictions. We show that every differentially private mechanism that handles insertions and deletions has worst-case additive error at least $T^{1/4}$ even under a relatively weak, event-level privacy definition. Then, we identify a parameter of the input stream, its maximum flippancy, that is low for natural data streams and for which we give tight parameterized error guarantees. Specifically, the maximum flippancy is the largest number of times that the contribution of a single item to the distinct elements count changes over the course of the stream. We present an item-level differentially private mechanism that, for all turnstile streams with maximum flippancy $w$, continually outputs the number of distinct elements with an $O(\sqrt{w}\cdot\mathsf{poly}\log T)$ additive error, without requiring prior knowledge of $w$. We prove that this is the best achievable error bound that depends only on $w$, for a large range of values of $w$. When $w$ is small, the error of our mechanism is similar to the polylogarithmic in $T$ error in the insertion-only setting, bypassing the hardness in the turnstile model.
@@ 1 Introduction
Machine learning algorithms are frequently run on sensitive data. In this context, a central challenge is to protect the privacy of individuals whose information is contained in the training set. Differential privacy [27] provides a rigorous framework for the design and analysis of algorithms that publish aggregate statistics, such as parameters of machine learning models, while preserving privacy. In this work, we focus on the model of differential privacy interchangeably called continual observation and continual release that was introduced by  Dwork et al. [25] and Chan et al. [13] to study privacy in settings when both the data and the published statics are constantly updated. One of the most fundamental statistics about a data stream is the number of distinct elements it contains (see, e.g., the book by Leskovec et al. [47]). The problem of counting distinct elements has been widely studied, starting with the work of Flajolet and Martin [32], and has numerous applications [2, 30, 49, 42, 4], including monitoring the number of logins from distinct accounts to a streaming service, tracking the number of different countries represented by people in a chat room, and tracking the number of students signed up for at least one club at a university. Algorithms for this problem are also used as basic building blocks in more complicated data analyses.
@@
Next, we define the function $\mathsf{CountDistinct}$ in the (turnstile) continual release model.
Definition 1.1 (Existence vector, $\mathsf{CountDistinct}$).
Fix a universe $\mathcal{U}$ and a time horizon $T\in\mathbb{N}$. For an element $u\in\mathcal{U}$ and a stream $x\in\mathcal{U}_{\pm}^{T}$, the existence vector $f_{u}(x)\in\{0,1\}^{T}$ is an indicator vector that tracks the existence of element $u$ in $x$: specifically, for each $t\in[T]$, the value $f_{u}(x)[t]$ is $1$ if and only if there are strictly more insertions than deletions of element $u$ in $x[1:t].$ The function $\mathsf{CountDistinct}{}:\mathcal{U}_{\pm}^{T}\to\mathbb{N}^{T}$ returns a vector of the same length as its input, where $\mathsf{CountDistinct}(x)[t]=\sum_{u\in\mathcal{U}}f_{u}(x)[t]$ for all $t\in[T]$.
The focus of our investigation is the best achievable error in the continual release model for a given time horizon $T$ and privacy parameters. We study the worst-case (over all input streams and time steps $t$) additive error of privately approximating the distinct elements counts under continual release.
Definition 1.2 (Error of an answer vector and error of a mechanism for $\mathsf{CountDistinct}$).
Given an answer vector $a\in\mathbb{R}^{T}$, the error of this vector with respect to the desired function value $f(x)\in\mathbb{R}^{T}$ computed on dataset $x$ is defined as $\text{\sf ERR}_{f}(x,a)=\|f(x)-a\|_{\infty}.$ A mechanism for $\mathsf{CountDistinct}$ in the continual release model is $\alpha$-accurate if it outputs a vector of answers $a$ with error $\text{\sf ERR}_{\mathsf{CountDistinct}}(x,a)\leq\alpha$ with probability at least 0.99.
Next, we discuss privacy. Originally, differential privacy [27] was defined in a setting where a data collector outputs the desired information about an entire dataset all at once. We call this the batch model to contrast it with continual release. In the batch model, two datasets are called neighbors if they differ in the data of one individual. There are two natural ways to adapt this definition to the continual release model [25, 13], depending on the desired privacy guarantees.
@@
Let $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$ be two streams of length $T$. Streams $x$ and $x^{\prime}$ are event-neighbors if one can be obtained from the other by replacing a stream entry with $\perp$. Streams $x$ and $x^{\prime}$ are item-neighbors if one can be obtained from the other by replacing a subset of stream entries pertaining to one specific element of $\mathcal{U}$ with symbols $\perp$.
Differential privacy can be defined with respect to any notion of neighboring datasets. There are two privacy parameters: $\varepsilon>0$ and $\delta\in[0,1)$. An algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private (DP) if for all pairs of neighboring datasets $x,x^{\prime}$ and all events $S$ in the output space of $\mathcal{A}$,
@@
The case when $\delta=0$ is referred to as pure differential privacy, and the general case as approximate differential privacy. For event-neighboring (respectively, item-neighboring) streams $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$, we say that $\mathcal{A}$ is $(\varepsilon,\delta)$-event-level-DP (respectively, item-level-DP). Item-level differential privacy imposes a more stringent requirement than event-level, since it guards against larger changes in the input stream. To contrast with the batch setting, we refer to continual release algorithms as mechanisms.
In the batch setting, where only $\mathsf{CountDistinct}(x)[T]$ is released, there is an $\varepsilon$-DP algorithm for counting distinct elements with expected error $O(1/\varepsilon)$ since the function $\mathsf{CountDistinct}(x)[T]$ has sensitivity 1—regardless of whether we consider event-level or item-level privacy. Privacy is more challenging in the continual release setting, where we aim to release a sequence of estimates, one for each time $t$, and we require that the privacy guarantee hold for the entire sequence of outputs. Prior work on privately estimating distinct elements in this setting considered the insertion-only model, exclusively: Bolot et al. [9] show that one can get a sequence of estimates, all of which are within additive error $\mathrm{poly}(\log T)/\varepsilon$. Their result holds for both item-level and event-level privacy (which are essentially equivalent for counting distinct elements with only insertions). Follow-up work generalized their mechanism but, again, considered only insertions [35, 29].
We uncover a much richer landscape in the turnstile model, even without considering memory restrictions. We show that any differentially private mechanism that handles insertions and deletions has worst-case additive error at least $T^{1/4}$ even under event-level privacy, the weaker of the two privacy notions. To overcome this lower bound, we identify a property of the input stream, its maximum flippancy, that is low for natural data streams and for which one can give tight parameterized error guarantees. To define flippancy, recall the notion of the existence vector from Definition 1.1.
@@
Given a stream $x$ of length $T$ and an element $u\in\mathcal{U}$, the flippancy of $u$ in $x$, denoted by $\mathsf{flip}(u,x)$, is the number of pairs of adjacent entries in the existence vector $f_{u}(x)$ with different values. That is, $\mathsf{flip}(u,x)=|\{j\in[T-1]:f_{u}(x)[j]\neq f_{u}(x)[j+1]\}|.$ The maximum flippancy of a stream $x$, denoted $w_{x}$, is $\max_{u\in\mathcal{U}}\mathsf{flip}(u,x)$.
In other words, the maximum flippancy is the largest number of times the contribution of a single item to the distinct elements count changes over the course of the stream. We design item-level private mechanisms whose error scales with the maximum flippancy of the stream, even though the maximum flippancy is not an input to the mechanism. We show matching lower bounds for item-level privacy that hold in all parameter regimes. For a large range of the flippancy parameter, we also show a matching lower bound for event-level privacy, via a different argument. This leaves a range with an intriguing gap between item-level and event-level bounds.
@@ 1.1 Our results
Bounds Event-Level Privacy Item-Level Privacy
@@
           $\tilde{O}\left(\mathsf{min}\left(\left(\sqrt{w_{x}}\log T+\log^{3}T\right)\cdot\frac{\sqrt{\log 1/\delta}}{\varepsilon},\frac{\left(T\log 1/\delta\right)^{1/3}}{\varepsilon^{2/3}},T\right)\right)$ (Thm. 1.5)
@@
$\Omega\left(\mathsf{min}\Big{(}\frac{\sqrt{w_{x}}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}},T\Big{)}\right)$ (Thm. 1.6) $\tilde{\Omega}\Big{(}\mathsf{min}\Big{(}\frac{\sqrt{w_{x}}}{\varepsilon},\frac{T^{1/3}}{\varepsilon^{2/3}},T\Big{)}\Big{)}$ (Thm. 1.7)
Table 1: Summary of our results: bounds on the worst-case additive error for $\mathsf{CountDistinct}$ under event-level and item-level $(\varepsilon,\delta)$-differential privacy, with $\varepsilon\leq 1$ and $\delta=o(\frac{\varepsilon}{T})$. The upper bound depends on the maximum flippancy $w_{x}$ of the input $x$, for every $x$. The lower bounds apply to the worst-case error of an algorithm taken over all inputs with a given maximum flippancy.
Our results are summarized in Table 1. Our first result is a mechanism for privately approximating $\mathsf{CountDistinct}$ for turnstile streams. For a stream $x$ of length $T$ with maximum flippancy $w_{x}$, this mechanism is item-level-DP and has error $\operatorname{O}\left(\min(\sqrt{w_{x}}\mdot\operatorname{polylog}T,T^{1/3}\right))$. Crucially, the mechanism is not given the maximum flippancy up front.
@@
For all $\varepsilon,\delta\in(0,1]$ and sufficiently large $T\in\mathbb{N}$, there exists an $(\varepsilon,\delta)$-item-level-DP mechanism for $\mathsf{CountDistinct}$ that is $\alpha$-accurate for all turnstile streams $x$ of length $T$, where
@@
Since this mechanism is item-level-DP, it is also event-level-DP with the same privacy parameters. The error it achieves is the best possible in terms of dependence only on $w_{x}$ for item-level privacy, and this error is nearly tight for event-level privacy. When $w_{x}$ is small, as is the case for many natural streams, our mechanism has error $O(\operatorname{polylog}T)$, similar to mechanisms for the insertion-only setting.
Theorem 1.5 can be easily extended to $\varepsilon$ bounded by any constant larger than $1$. We fixed the bound to be $1$ to simplify the presentation. Our mechanism has polynomial time and space complexity in the input parameters, although it does not achieve the typically sublinear space guarantees of streaming algorithms. (See “Bounded Memory” in Section 1.4 for discussion.)
Our lower bounds on the accuracy for $\mathsf{CountDistinct}$ for turnstile streams are parametrized by a flippancy bound $w$, and apply for streams with maximum flippancy $w_{x}\leq w$. For event-level DP, our lower bound shows that for all mechanisms with error guarantees expressed solely in terms of the maximum flippancy $w_{x}$, time horizon $T$, and privacy parameter $\varepsilon$, our $\mathsf{CountDistinct}$ mechanism is asymptotically optimal for a large range of values of $w_{x}$, namely, for all $w_{x}\leq T^{1/2}$ and $w_{x}\geq T^{2/3}$. The best achievable error for $w_{x}\in(T^{1/2},T^{2/3})$ for event-level differential privacy remains an open problem.
@@
For all $\varepsilon,\delta\in(0,1]$, sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$, and all $(\varepsilon,\delta)$-event-level-DP mechanisms that are $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, if $\delta=o(\frac{\varepsilon}{T})$, then
@@
In particular, for every $w$, there exists a family of streams satisfying $w_{x}=\Theta(w)$, for which every mechanism has worst-case error $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w_{x}}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}},T\right)\right)$. This is reflected in Table 1.
For item-level-DP, our lower bound on the error matches our upper bound for all regimes of $w_{x}$ up to polylogarithmic factors.
@@
Let $\varepsilon\in(0,1]$, $\delta\in(0,1]$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. For all $(\varepsilon,\delta)$-item-level-DP mechanisms that are $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$:
1 If $\delta=o(\varepsilon/T)$, then $\alpha=\tilde{\Omega}\Big{(}\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/3}}{\varepsilon^{2/3}},T)\Big{)}\Big{)}$.
2 If $\delta=0$, then $\alpha=\Omega\Big{(}\mathsf{min}\Big{(}\frac{w}{\varepsilon},\sqrt{\frac{T}{\varepsilon}},T\Big{)}\Big{)}$.
Like for Theorem 1.6, we can replace $w$ with $w_{x}$ in the lower bounds stated in Theorem 1.7 in the following sense: for every $w$, there exists a family of streams satisfying $w_{x}=\Theta(w)$, for which every mechanism has worst-case error as stated, but with $w$ replaced with $w_{x}$. (This applies to both Item 1 and Item 2.) For Item 1, this is reflected in Table 1.
Variants of the model.
All our lower bounds also hold in the strict turnstile model, where element counts never go below 00. They also apply to offline mechanisms that receive the entire input stream before producing output; they do not rely on the mechanism’s uncertainty about what comes later in the stream. Furthermore, our item-level lower bounds hold even in the model where each element can be inserted only when it is absent and deleted only when it is present. We call this the likes model, since it captures the behavior of the “like” counts on social media websites. Our event-level lower bound does not apply to the likes model, which admits an event-level-DP mechanism for $\mathsf{CountDistinct}$ with error polylogarithic in $T$, as explained in Section 1.3 in the discussion of the work of Henzinger et al. [40]. In Appendix B, we provide formal statements and proof outlines for the aforementioned lower bounds for the variants of our model.